There were 600 kids invited to an ice cream party,7/10 of the kids had vanilla ice cream,1/4 had chocolate ice cream.The rest had an ice cream sandwich,
How many kids had a vanilla ice cream cone and how many kids had an ice cream sandwich?

Answer by checkley77(12569)

.7*600=420 kids had vanilla ice cream.
.25*600=150 kids had chocolate ice cream.
600-420-150=600-570=30 kids had an ice cream sandwich..